OpenStudy (anonymous):

Solve the following system of equations algebraically:? Y=400+20X, Y=600+10X

OpenStudy (anonymous):

well y=y so you can just set them = to each other 400+20x=600+10x solve for x

OpenStudy (anonymous):

Each of these equations represents lines on a graph, the "solution" is the point where the two lines cross, thus to solve, you set the equations as equal to one another.

OpenStudy (anonymous):

You can check your answer with a graphing calculator

OpenStudy (anonymous):

I NEED HELP WITH THE STEPS.

OpenStudy (anonymous):

well.. 400+20x=600+10x -200=10x ... finish solving for x then plug into one of the equations for y

OpenStudy (anonymous):

-20x+y=400 -10x+y=600 We can use elimination for this. Multiply 2nd equation by -2 and then add together to get rid of x. -20x+y=400 20x-2y=-1200 =>-y=-800 y=800 Then plug in y and solve for x. -20x+800=400 =>-20x=-400 x=20 Try it in the other equation. -10x+400=600 -10x=-200 x=20 Hope this helps. Thanks, Sean Walsh Facebook: Tutor Sean http://www.tutorsean.net

OpenStudy (anonymous):

WHERE DID YOU GET THE -20 AND THE -10 IN THE EQUATIONS? IT'S A POSITIVE 10 AND 20?

OpenStudy (anonymous):

Using algebra to manipulate across the equals sign. Y=400+20X is the same as Y-20X=400+20X-20X. :)

OpenStudy (anonymous):

Did that help clarify? You can use the gaussian elimination method as shown above to solve these types of equations and it makes it really easy. This is a method which is taught in some algebra classes but mainly in Linear Algebra, which shows numerous techniques. If you check out my site, there are a few free videos on linear algebra shortcuts you may find useful for systems of equations. http://www.tutorsean.net

OpenStudy (anonymous):

algebraically determines when the two savings plans yield identical balances and accurately states the balance. ...using the two equation this is what i need to find....Solve the system of equations algebraically to determine when the two savings plans yield identical balances and state this balance.

OpenStudy (anonymous):

You will find the same answer this way. Here is what I think you are talking about. 400+20x=600+10x 20x=200+10x 10x=200 x=20 Now plug in for x in either equation to find y.

OpenStudy (anonymous):

confused...so i could just plug in any #?